OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 2x − 3y = 3 and passes through the point (−8, 2). y = x − 10 y = x + 10 y = −x − 10 y = −x + 10 please help!

jimthompson5910 (jim_thompson5910):

2x − 3y = 3 -3y = -2x + 3 -3y/(-3) = -2x/(-3) + 3/(-3) y = 2/3x - 3

jimthompson5910 (jim_thompson5910):

The slope here is 2/3 flip the fraction and flip the sign to get -3/2

jimthompson5910 (jim_thompson5910):

y = mx + b y = (-3/2)x + b 2 = (-3/2)*(-8) + b 2 = 24/2 + b 2 = 12 + b 2 - 12 = b -10 = b b = -10 So the equation of the perpendicular line is y = -3/2x - 10

OpenStudy (anonymous):

To check your answer,you can simply substitute the values of x and y in both the cases and see in which case LHS=RHS and that will the answer

OpenStudy (campbell_st):

well I think you have a typo in your question ans none of the answer choices match the correct solution.

jimthompson5910 (jim_thompson5910):

the fractions didn't copy/paste correctly campbell_st

OpenStudy (anonymous):

why 2nd one is a match

OpenStudy (anonymous):

2=-8+10 2=2